OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (ivettef365):

convert the equation as y=mx+b where m=slope and b=y-int

OpenStudy (anonymous):

what.? if i knew how to do that i wouldnt be in summer school.

OpenStudy (anonymous):

y=mx+b -6=5x-3y Sooo 5=slope and 3=y-intercept

OpenStudy (anonymous):

putting x=0 we get the y int so 5(0)-3y=-6 so y=2 as y int is the point on y axis so that points x component is 0 and the slope can be found out by differentiating the equation and finding dy/dx

OpenStudy (anonymous):

i am not positive though

OpenStudy (ivettef365):

to isolate the y first you move the 5x, when you move to the other side change the sign -3y = -5x -6

OpenStudy (ivettef365):

then divide by -3 you have y = 5/3x+2

OpenStudy (anonymous):

then divide by -3

OpenStudy (ivettef365):

now you have it in the right format, 5/3 = m = slope and 2 = b = y-int